Flask 'A' of volume 10 L containing 20g of H2 and flask 'B' of volume
10 L containing 88 g of CO2 are connected by a connector having
negligible volume. When the valve of the connector is opened, the
number of moles of H2 remains in the flask 'A' is
T(H2)=500k 
T(CO2)=200K

Grade:11

1 Answers

Shweta Ahire
5 years ago
Initial moles of H2 in flask A = 10
Initial moles of CO2 in flask B = 2 
Total moles = 12 
After opening the connector ,  gases will mix then pressure of both of the flask changes but no of total moles be same
So p×10/R× 500  +  p×10/R × 200 = 12(total moles)
After solving this, p= 1200R/ 7
moles remaining in flask A =(1200R/7 × 10)/R ×500
= 24/7  =  3.428